OpenStudy (anonymous):

9x^2(3x^34y) simplify

Parth (parthkohli):

Don't let the parentheses intimidate you.

OpenStudy (anonymous):

lol

Parth (parthkohli):

\[9x^2(3x^34y) \implies 9x^2 \times 3x^3\times 4y \]

Parth (parthkohli):

And just combine those coefficients and \(x\)'s.

OpenStudy (anonymous):

9x^2(3x^3)9x^2(4y)=27x^51*36x^2y^4

OpenStudy (anonymous):

thanks